# Increasing Batch Size Improves Convergence of Stochastic Gradient Descent with Momentum

### v1

@@ Abstract
Stochastic gradient descent with momentum (SGDM), which is defined by adding a momentum term to SGD, has been well studied in both theory and practice. Theoretically investigated results showed that the settings of the learning rate and momentum weight affect the convergence of SGDM. Meanwhile, practical results showed that the setting of batch size strongly depends on the performance of SGDM. In this paper, we focus on mini-batch SGDM with constant learning rate and constant momentum weight, which is frequently used to train deep neural networks in practice. The contribution of this paper is showing theoretically that using a constant batch size does not always minimize the expectation of the full gradient norm of the empirical loss in training a deep neural network, whereas using an increasing batch size definitely minimizes it, that is, increasing batch size improves convergence of mini-batch SGDM. We also provide numerical results supporting our analyses, indicating specifically that mini-batch SGDM with an increasing batch size converges to stationary points faster than with a constant batch size. Python implementations of the optimizers used in the numerical experiments are available at https://anonymous.4open.science/r/momentum-increasing-batch-size- [^8] C/.
Machine Learning, ICML
@@ 1 Introduction
Stochastic gradient descent (SGD) and its variants, such as SGD with momentum (SGDM) and adaptive methods, are well known as useful optimizers for minimizing the empirical loss defined by the mean of nonconvex loss functions in training a deep neural network (DNN). In the present paper, we focus on SGDM optimizers, which are defined by adding a momentum term to SGD. Various types of SGDM have been proposed, such as stochastic heavy ball (SHB) (Polyak, 1964), normalized-SHB (NSHB) (Gupal & Bazhenov, 1972), Nesterov’s accelerated gradient method (Nesterov, 1983; Sutskever et al., 2013), synthesized Nesterov variants (Lessard et al., 2016), Triple Momentum (Van Scoy et al., 2018), Robust Momentum (Cyrus et al., 2018), PID control-based methods (An et al., 2018), stochastic unified momentum (SUM) (Yan et al., 2018), accelerated SGD (Jain et al., 2018; Kidambi et al., 2018; Varre & Flammarion, 2022; Li et al., 2024), quasi-hyperbolic momentum (QHM) (Ma & Yarats, 2019), and proximal-type SHB (PSHB) (Mai & Johansson, 2020).
Since the empirical loss is nonconvex with respect to a parameter $\bm{\theta}\in\mathbb{R}^{d}$ of a DNN, we are interested in nonconvex optimization for SGDM. Let $\bm{\theta}_{t}\in\mathbb{R}^{d}$ be the $t$-th approximation of SGDM to minimize the nonconvex empirical loss function $f\colon\mathbb{R}^{d}\to\mathbb{R}$. SGDM is defined as $\bm{\theta}_{t+1}=\bm{\theta}_{t}-\eta_{t}\bm{m}_{t}$, where $\eta_{t}>0$ is a learning rate and $\bm{m}_{t}$ is a momentum buffer. For example, SGDM with $\bm{m}_{t}:=\beta\bm{m}_{t-1}+\nabla f_{B_{t}}(\bm{\theta}_{t})$ is SHB, where $\nabla f_{B_{t}}\colon\mathbb{R}^{d}\to\mathbb{R}^{d}$ denotes the stochastic gradient of $f$ and $\beta\in[0,1)$ is a momentum weight. SGDM with $\bm{m}_{t}:=\beta\bm{m}_{t-1}+(1-\beta)\nabla f_{B_{t}}(\bm{\theta}_{t})$ is NSHB. Since SHB with $\beta=0$ (NSHB with $\beta=0$) coincides with SGD defined by $\bm{\theta}_{t+1}=\bm{\theta}_{t}-\eta_{t}\nabla f_{B_{t}}(\bm{\theta}_{t})$, SGDM is defined by adding a momentum term (e.g., $\beta\bm{m}_{t-1}$ in the case of SHB) to SGD.
[…]
@@
[Theorem 3.2] Mini-batch
@@
[Theorem 3.4] Mini-batch
Table 1 summarizes convergence analyses of SGDM for nonconvex optimization. For example, NSHB ((6) in Table 1) using a constant learning rate $\eta_{t}=\eta>0$ and a constant momentum weight $\beta_{t}=\beta$ satisfies $\min_{t\in[0:T-1]}\mathbb{E}[\|\nabla f(\bm{\theta}_{t})\|]=O(\sqrt{\frac{1}{T}+\sigma^{2}})$ (Liu et al., 2020, Theorem 1), where $T$ is the number of steps, $\nabla f\colon\mathbb{R}^{d}\to\mathbb{R}^{d}$ is the gradient of $f$, $\sigma^{2}$ is the upper bound of the variance of the stochastic gradient of $f$, and $\mathbb{E}[X]$ denotes the expectation of a random variable $X$. In comparison, QHM ((2) in Table 1), which is a generalization of NSHB, using a decaying learning rate $\eta_{t}$ and a decaying momentum weight $\beta_{t}$ satisfies $\liminf_{t\to+\infty}\|\nabla f(\bm{\theta}_{t})\|=0$ (Gitman et al., 2019, Theorem 1). As can be seen from these convergence analysis results, the performance of SGDM in finding a stationary point $\bm{\theta}^{\star}$ of $f$ (i.e., $\nabla f(\bm{\theta}^{\star})=\bm{0}$) depends on the settings of the learning rate $\eta_{t}$ and the momentum weight $\beta_{t}$.
Moreover, we would like to emphasize that the setting of the batch size $b_{t}$ affects the performance of SGDM. Previous results in (Shallue et al., 2019; Zhang et al., 2019) numerically showed that, for deep learning optimizers, the number of steps needed to train a DNN halves for each doubling of the batch size. In (Smith et al., 2018), it was numerically shown that using an enormous batch size leads to a reduction in the number of parameter updates and the model training time. Hence, we decided to investigate theoretically how the setting of batch size affects convergence of SGDM.
@@ 1.1 Contribution
In this paper, we focus on mini-batch SGDM with constant learning rate $\eta>0$ and constant momentum weight $\beta\in[0,1)$, which is frequently used to train DNNs in practice.
@@
    which implies that mini-batch SGDM does not always minimize the expectation of the full gradient norm of the empirical loss in training a DNN (Table 1; Theorems 3.1 and 3.2).
The bias term $\frac{f(\bm{\theta}_{0})-f^{\star}}{\eta T}$ converges to 00 when $T\to+\infty$. However, the variance term $\frac{L\eta\sigma^{2}}{b}$ remains a constant positive real number regardless of how large $T$ is. In contrast, using a large batch size $b$ makes the variance term $\frac{L\eta\sigma^{2}}{b}$ small. Hence, we can expect that an upper bound of $\min_{t\in[0:T-1]}\mathbb{E}[\|\nabla f(\bm{\theta}_{t})\|]$ for mini-batch SGDM with increasing batch size converges to 00.
@@
    which implies that mini-batch SGDM minimizes the expectation of the full gradient norm of the empirical loss in the sense of an $O(\frac{1}{\sqrt{T}})$ rate of convergence (Table 1; Theorems 3.3 and 3.4).
The previous results reported in (Byrd et al., 2012; Balles et al., 2016; De et al., 2017; Smith et al., 2018; Goyal et al., 2018; Shallue et al., 2019; Zhang et al., 2019) indicated that increasing batch sizes is useful for training DNNs with deep learning optimizers. However, providing the theoretical performance of mini-batch SGDM with an increasing batch size may be insufficient, as seen in the existing analyses of SGDM (Table 1). The paper shows theoretically that SGDM with an increasing batch size converges to stationary points of the empirical loss (Theorems 3.3 and 3.4). The previous results in (Yan et al., 2018, Theorem 1), (Mai & Johansson, 2020, Theorem 1), and (Yu et al., 2019, Corollary 1) (Table 1(1), (4), and (5)) showed that SGDM with constant learning rate $\eta=O(\frac{1}{\sqrt{T}})$ and a constant momentum weight $\beta$ has convergence rate $O(\frac{1}{T^{1/4}})$. Our results (Theorems 3.3 and 3.4) guarantee that, if the batch size increases, then SGDM satisfies $\min_{t\in[0:T-1]}\mathbb{E}[\|\nabla f(\bm{\theta}_{t})\|]=O(\frac{1}{T^{1/2}})$, which improves the previous convergence rate $O(\frac{1}{T^{1/4}})$.
The result in (1) indicates that the performance of mini-batch SGDM with increasing batch size $b_{t}$ depends on $\delta$. Let $\eta$ and $\beta$ be fixed (e.g., $\eta=0.1$ and $\beta=0.9$). Then, (1) indicates that the larger $\delta$ is, the smaller the variance term $\frac{L\eta\sigma^{2}\delta}{(\beta^{2}\delta-1)b_{0}T}$ is (since $\frac{\delta}{\beta^{2}\delta-1}=\frac{1}{(0.9)^{2}-1/\delta}$ becomes small as $\delta$ becomes large). We are interested in verifying whether this theoretical result holds in practice.
@@
    The third contribution is showing numerically that quadruply increasing batch size (i.e., $\delta=4$) decreases $\min_{t\in[0:T-1]}\|\nabla f(\bm{\theta}_{t})\|$ faster than doubly increasing batch size (i.e., $\delta=2$) or maintaining a constant batch size.
We consider training ResNet-18 on the CIFAR-100 and Tiny ImageNet datasets using not only NSHB and SHB but also baseline optimizers: SGD, Adam (Kingma & Ba, 2015), AdamW (Loshchilov & Hutter, 2019), and RMSprop (Tieleman & Hinton, 2012). A particularly interesting result in Section 4 is that an increasing batch size is applicable for Adam in the sense of minimizing $\min_{t\in[0:T-1]}\|\nabla f(\bm{\theta}_{t})\|$ fastest. Hence, in the future, we would like to verify whether Adam with an increasing batch size theoretically has a better convergence rate than SGDM.
@@ 2.1 Empirical risk minimization
[…]
We assume that the loss functions $f_{i}$ ($i\in[n]$) satisfy the conditions in the following assumption.
Assumption 2.1.
@@
(A2) Let $\xi$ be a random variable independent of $\bm{\theta}\in\mathbb{R}^{d}$. $\nabla f_{\xi}\colon\mathbb{R}^{d}\to\mathbb{R}^{d}$ is the stochastic gradient of $\nabla f$ such that (i) for all $\bm{\theta}\in\mathbb{R}^{d}$, $\mathbb{E}_{\xi}[\nabla f_{\xi}(\bm{\theta})]=\nabla f(\bm{\theta})$ and (ii) there exists $\sigma\geq 0$ such that, for all $\bm{\theta}\in\mathbb{R}^{d}$, $\mathbb{V}_{\xi}[\nabla f_{\xi}(\bm{\theta})]=\mathbb{E}_{\xi}[\|\nabla f_{\xi}(\bm{\theta})-\nabla f(\bm{\theta})\|^{2}]\leq\sigma^{2}$, where $\mathbb{E}_{\xi}[\cdot]$ denotes expectation with respect to $\xi$.
(A3) Let $b\in\mathbb{N}$ such that $b\leq n$ and let $\bm{\xi}=(\xi_{1},\xi_{2},\cdots,\xi_{b})^{\top}$ comprise $b$ independent and identically distributed variables and be independent of $\bm{\theta}\in\mathbb{R}^{d}$. The full gradient $\nabla f(\bm{\theta})$ is estimated as the mini-batch gradient at $\bm{\theta}$ defined by $\nabla f_{B}(\bm{\theta}):=\frac{1}{b}\sum_{i=1}^{b}\nabla f_{\xi_{i}}(\bm{\theta})$.

### v2

@@ Abstract
Stochastic gradient descent with momentum (SGDM), in which a momentum term is added to SGD, has been well studied in both theory and practice. The theoretical studies show that the settings of the learning rate and momentum weight affect the convergence of SGDM. Meanwhile, the practical studies have shown that the batch-size setting strongly affects the performance of SGDM. In this paper, we focus on mini-batch SGDM with a constant learning rate and constant momentum weight, which is frequently used to train deep neural networks. We show theoretically that using a constant batch size does not always minimize the expectation of the full gradient norm of the empirical loss in training a deep neural network, whereas using an increasing batch size definitely minimizes it; that is, an increasing batch size improves the convergence of mini-batch SGDM. We also provide numerical results supporting our analyses, indicating specifically that mini-batch SGDM with an increasing batch size converges to stationary points faster than with a constant batch size, while also reducing computational cost. Python implementations of the optimizers used in the numerical experiments are available at https://github.com/iiduka-researches/NSHB_increasing_batchsize_acml25/.
keywords:
convergence rate; increasing batch size; nonconvex optimization; stochastic gradient descent with momentum
editors: Hung-yi Lee and Tongliang Liu
@@ 1 Introduction
Stochastic gradient descent (SGD) and its variants, such as SGD with momentum (SGDM) and adaptive methods, are useful optimizers for minimizing the empirical loss defined by the mean of nonconvex loss functions in training a deep neural network (DNN). In the present paper, we focus on SGDM optimizers, in which a momentum term is added to SGD. Various types of SGDM have been proposed, including stochastic heavy ball (SHB) (Polyak, 1964), normalized-SHB (NSHB) (Gupal and Bazhenov, 1972), Nesterov’s accelerated gradient method (Nesterov, 1983; Sutskever et al., 2013), synthesized Nesterov variants (Lessard et al., 2016), Triple Momentum (Van Scoy et al., 2018), Robust Momentum (Cyrus et al., 2018), PID control-based methods (An et al., 2018), stochastic unified momentum (SUM) (Yan et al., 2018), accelerated SGD (Jain et al., 2018; Kidambi et al., 2018; Varre and Flammarion, 2022; Li et al., 2024), quasi-hyperbolic momentum (QHM) (Ma and Yarats, 2019), and proximal-type SHB (PSHB) (Mai and Johansson, 2020).
Since the empirical loss is nonconvex with respect to a parameter $\bm{\theta}\in\mathbb{R}^{d}$ of a DNN, we are interested in nonconvex optimization for SGDM. Let $\bm{\theta}_{t}\in\mathbb{R}^{d}$ be the $t$-th approximation of SGDM to minimize the nonconvex empirical loss function $f\colon\mathbb{R}^{d}\to\mathbb{R}$. SGDM is defined as $\bm{\theta}_{t+1}=\bm{\theta}_{t}-\eta_{t}\bm{m}_{t}$, where $\eta_{t}>0$ is a learning rate and $\bm{m}_{t}$ is a momentum buffer. For example, SGDM with $\bm{m}_{t}:=\beta\bm{m}_{t-1}+\nabla f_{B_{t}}(\bm{\theta}_{t})$ is SHB, where $\nabla f_{B_{t}}\colon\mathbb{R}^{d}\to\mathbb{R}^{d}$ denotes the stochastic gradient of $f$ and $\beta\in[0,1)$ is a momentum weight. SGDM with $\bm{m}_{t}:=\beta\bm{m}_{t-1}+(1-\beta)\nabla f_{B_{t}}(\bm{\theta}_{t})$ is NSHB. Since SHB with $\beta=0$ (NSHB with $\beta=0$) coincides with SGD, which defined by $\bm{\theta}_{t+1}=\bm{\theta}_{t}-\eta_{t}\nabla f_{B_{t}}(\bm{\theta}_{t})$, SGDM is defined as SGD with an added momentum term (e.g., $\beta\bm{m}_{t-1}$ in the case of SHB).
Table 1: Convergence of SGDM optimizers to minimize $L$-smooth $f$ over the number of steps $T$. “Noise” in the Gradient column means that the optimizer uses noisy observations, i.e., $\bm{g}(\bm{\theta})=\nabla f(\bm{\theta})+\text{(Noise)}$, of the full gradient $\nabla f(\bm{\theta})$, where $\sigma^{2}$ is an upper bound of Noise, while “Increasing (resp. Constant) Mini-batch” in the Gradient column means that the optimizer uses a mini-batch gradient $\nabla f_{B_{t}}(\bm{\theta})=\frac{1}{b_{t}}\sum_{i=1}^{b_{t}}\nabla f_{\xi_{t,i}}(\bm{\theta})$ with a batch size $b_{t}$ such that $b_{t}\leq b_{t+1}$ (resp. $b_{t}=b$). “Bounded Gradient” in the Additional Assumption column means that there exists $G>0$ such that, for all $t\in\mathbb{N}$, $\|\nabla f(\bm{\theta}_{t})\|\leq G$, where $(\bm{\theta}_{t})_{t=0}^{T-1}$ is the sequence generated by the optimizer. “Polyak-Łojasiewicz” in the Additional Assumption column means that there exists $\rho>0$ such that, for all $t\in\mathbb{N}$, $\|\nabla f(\bm{\theta}_{t})\|^{2}\geq 2\rho(f(\bm{\theta}_{t})-f^{\star})$, where $f^{\star}$ is the optimal value of $f$ over $\mathbb{R}^{d}$. Here, we let $\mathbb{E}\|\nabla f_{T}\|:=\min_{t\in[0:T-1]}\mathbb{E}[\|\nabla f(\bm{\theta}_{t})\|]$. Results (1)–(7) were presented in (1) (Yan et al., 2018, Theorem 1), (2) (Gitman et al., 2019, Theorem 1), (3) (Gitman et al., 2019, Theorem 2), (4) (Mai and Johansson, 2020, Theorem 1), (5) (Yu et al., 2019, Corollary 1), (6) (Liu et al., 2020, Theorem 1), and (7) (Liang et al., 2023, Theorem 4.1).
@@
[Theorem A.1]] Mini-batch
@@
[Theorem A.2]] Mini-batch
Table 1 summarizes the convergence analyses of SGDM for nonconvex optimization. For example, NSHB ((6) in Table 1) using a constant learning rate $\eta_{t}=\eta>0$ and a constant momentum weight $\beta_{t}=\beta$ satisfies $\min_{t\in[0:T-1]}\mathbb{E}[\|\nabla f(\bm{\theta}_{t})\|]=O(\sqrt{\frac{1}{T}+\sigma^{2}})$ (Liu et al., 2020, Theorem 1), where $T$ is the number of steps, $\nabla f\colon\mathbb{R}^{d}\to\mathbb{R}^{d}$ is the gradient of $f$, $\sigma^{2}$ is the upper bound of the variance of the stochastic gradient of $f$, and $\mathbb{E}[X]$ denotes the expectation of a random variable $X$. In comparison, QHM ((2) in Table 1), which is a generalization of NSHB, using a decaying learning rate $\eta_{t}$ and a decaying momentum weight $\beta_{t}$ satisfies $\liminf_{t\to+\infty}\|\nabla f(\bm{\theta}_{t})\|=0$ (Gitman et al., 2019, Theorem 1). As can be seen from these convergence analysis results, the performance of SGDM in finding a stationary point $\bm{\theta}^{\star}$ of $f$ (i.e., $\nabla f(\bm{\theta}^{\star})=\bm{0}$) depends on the settings of the learning rate $\eta_{t}$ and the momentum weight $\beta_{t}$.
Moreover, we would like to emphasize that the setting of the batch size $b_{t}$ affects the performance of SGDM. Previous results presented in (Shallue et al., 2019; Zhang et al., 2019) numerically showed that, for deep learning optimizers, the number of steps needed to train a DNN is halved for each doubling of the batch size. In (Smith et al., 2018), it was numerically shown that using an enormous batch size leads to a reduction in the number of parameter updates and the model training time. In addition, related research has explored plain SGD, highlighting the broader interest in batch-size strategies beyond SGDM. For instance, the two-scale adaptive (TSA) method (Gao et al., 2022) co-adapts the batch size and step size to obtain exact convergence and favorable sample complexity. The related work has also investigated batch-size growth for Riemannian SGD (Sakai and Iiduka, 2025). While these studies share the idea of using an increasing batch size, their analyses are limited to plain SGD. Therefore, in this paper, we theoretically investigate how the setting of the batch size affects the convergence of SGDM under constant hyperparameters.
@@ 1.1 Contribution
In this paper, we focus on mini-batch SGDM with a constant learning rate $\eta>0$ and constant momentum weight $\beta\in[0,1)$, which is frequently used to train DNNs.
@@
    which implies that mini-batch SGDM does not always minimize the expectation of the full gradient norm of the empirical loss in training a DNN (Table 1; Theorems 3.1 and A.1).
The bias term $\frac{f(\bm{\theta}_{0})-f^{\star}}{\eta T}$ converges to 0 when $T\to+\infty$. However, the variance term $\frac{L\eta\sigma^{2}}{b}$ remains a constant positive real number regardless of how large $T$ is. In contrast, using a large batch size $b$ makes the variance term $\frac{L\eta\sigma^{2}}{b}$ small. Hence, we can expect that the upper bound of $\min_{t\in[0:T-1]}\mathbb{E}[\|\nabla f(\bm{\theta}_{t})\|]$ for mini-batch SGDM with increasing batch size converges to 0.
@@
    which implies that mini-batch SGDM minimizes the expectation of the full gradient norm of the empirical loss in the sense of an $O(\frac{1}{\sqrt{T}})$ rate of convergence (Table 1; Theorems 3.2 and A.2).
The previous results reported in (Byrd et al., 2012; Balles et al., 2016; De et al., 2017; Smith et al., 2018; Goyal et al., 2018; Shallue et al., 2019; Zhang et al., 2019) indicated that increasing batch sizes are useful for training DNNs with deep-learning optimizers. However, the existing analyses of SGDM have indicated that knowing only the theoretical performance of mini-batch SGDM with an increasing batch size may be insufficient (Table 1). The paper shows theoretically that SGDM with an increasing batch size converges to stationary points of the empirical loss (Theorems 3.2 and A.2). The previous results in (Yan et al., 2018, Theorem 1), (Mai and Johansson, 2020, Theorem 1), and (Yu et al., 2019, Corollary 1) (Table 1(1), (4), and (5)) showed that SGDM with a constant learning rate $\eta=O(\frac{1}{\sqrt{T}})$ and a constant momentum weight $\beta$ has convergence rate $O(\frac{1}{T^{1/4}})$. Our results (Theorems 3.2 and A.2) guarantee that, if the batch size increases, then SGDM satisfies $\min_{t\in[0:T-1]}\mathbb{E}[\|\nabla f(\bm{\theta}_{t})\|]=O(\frac{1}{T^{1/2}})$, which is an improvement on the previous convergence rate $O(\frac{1}{T^{1/4}})$.
The result in (1) indicates that the performance of mini-batch SGDM with an increasing batch size $b_{t}$ depends on $\delta$. Let $\eta$ and $\beta$ be fixed (e.g., $\eta=0.1$ and $\beta=0.9$). Then, (1) indicates that the larger $\delta$ is, the smaller the variance term $\frac{L\eta\sigma^{2}\delta}{(\beta^{2}\delta-1)b_{0}T}$ becomes (since $\frac{\delta}{\beta^{2}\delta-1}=\frac{1}{(0.9)^{2}-1/\delta}$ becomes small as $\delta$ becomes large). We are interested in verifying whether this theoretical result holds in practice. Our numerical results in Section 4 support the theoretical findings, suggesting that an increasing batch size leads to faster convergence.
We trained ResNet-18 on the CIFAR-100 and Tiny-ImageNet datasets by using not only NSHB and SHB but also baseline optimizers: SGD, Adam (Kingma and Ba, 2015), AdamW (Loshchilov and Hutter, 2019), and RMSprop (Tieleman and Hinton, 2012). The results on Tiny-ImageNet are provided in Appendix A.4. A particularly interesting result in Section 4 is that an increasing batch size benefits Adam in the sense of minimizing $\min_{t\in[0:T-1]}\|\nabla f(\bm{\theta}_{t})\|$ fastest. Hence, in the future, we would like to verify whether Adam with an increasing batch size theoretically has a better convergence rate than SGDM.
@@
    The third contribution of this paper is to show that an increasing batch size significantly reduces the computational cost required to achieve optimization and generalization criteria.
In this paper, we define the computational cost as the stochastic first-order oracle (SFO) complexity, in accordance with prior work that formalized this concept (Sato and Iiduka, 2023; Imaizumi and Iiduka, 2024), where SFO is interpreted as the total number of stochastic gradient evaluations. This perspective resonates with the extensive empirical investigation made by Shallue et al. (Shallue et al., 2019), which systematically examined how the batch size affects training efficiency across diverse neural network settings. Let $T$ be the number of training steps and $b$ the batch size. Then, the total SFO complexity is given by $Tb$ when using a fixed batch size. Our numerical results in Section 4.2 demonstrate that, under realistic GPU memory constraints, SGDM with an increasing batch size requires significantly fewer gradient computations to achieve competitive optimization and generalization performance compared with using a fixed batch size. These results indicate that an increasing batch size is not only theoretically appealing but also practically effective in reducing the computational burden in deep-learning training.
@@ 2.1 Empirical risk minimization
Let $\bm{\theta}\in\mathbb{R}^{d}$ be a parameter of a DNN, where $\mathbb{R}^{d}$ is $d$-dimensional Euclidean space with inner product $\langle\cdot,\cdot\rangle$ and induced norm $\|\cdot\|$. Let $\mathbb{R}_{+}:=\{x\in\mathbb{R}\colon x\geq 0\}$ and $\mathbb{R}_{++}:=\{x\in\mathbb{R}\colon x>0\}$. Let $\mathbb{N}$ be the set of natural numbers. Let $S=\{(\bm{x}_{1},\bm{y}_{1}),\ldots,(\bm{x}_{n},\bm{y}_{n})\}$ be the training set, where the data point $\bm{x}_{i}$ is associated with a label $\bm{y}_{i}$ and $n\in\mathbb{N}$ is the number of training samples. Let $f_{i}(\cdot):=f(\cdot;(\bm{x}_{i},\bm{y}_{i}))\colon\mathbb{R}^{d}\to\mathbb{R}_{+}$ be the loss function corresponding to the $i$-th labeled training data $(\bm{x}_{i},\bm{y}_{i})$. Empirical risk minimization (ERM) minimizes the empirical loss defined for all $\bm{\theta}\in\mathbb{R}^{d}$ as $f(\bm{\theta})=\frac{1}{n}\sum_{i\in[n]}f(\bm{\theta};(\bm{x}_{i},\bm{y}_{i}))=\frac{1}{n}\sum_{i\in[n]}f_{i}(\bm{\theta})$, where $n]:=\{1,2,\cdots,n\}$.
We assume that the loss functions $f_{i}$ ($i\in[n]$) satisfy the following conditions.
Assumption 1.
@@
(A2) Let $\xi$ be a random variable independent of $\bm{\theta}\in\mathbb{R}^{d}$. $\nabla f_{\xi}\colon\mathbb{R}^{d}\to\mathbb{R}^{d}$ is the stochastic gradient of $\nabla f$ such that (i) for all $\bm{\theta}\in\mathbb{R}^{d}$, $\mathbb{E}_{\xi}[\nabla f_{\xi}(\bm{\theta})]=\nabla f(\bm{\theta})$ and (ii) there exists $\sigma\geq 0$ such that, for all $\bm{\theta}\in\mathbb{R}^{d}$, $\mathbb{V}_{\xi}[\nabla f_{\xi}(\bm{\theta})]=\mathbb{E}_{\xi}[\|\nabla f_{\xi}(\bm{\theta})-\nabla f(\bm{\theta})\|^{2}]\leq\sigma^{2}$, where $\mathbb{E}_{\xi}[\cdot]$ denotes the expectation with respect to $\xi$.
(A3) Let $b\in\mathbb{N}$ such that $b\leq n$ and let $\bm{\xi}=(\xi_{1},\xi_{2},\cdots,\xi_{b})^{\top}$ comprise $b$ independent and identically distributed variables that are independent of $\bm{\theta}\in\mathbb{R}^{d}$. The full gradient $\nabla f(\bm{\theta})$ is taken to be the mini-batch gradient at $\bm{\theta}$ defined by $\nabla f_{B}(\bm{\theta}):=\frac{1}{b}\sum_{i=1}^{b}\nabla f_{\xi_{i}}(\bm{\theta})$.